OpenStudy (anonymous):

what is the multiplicative of -2 3/4

OpenStudy (jdoe0001):

the w0t?

OpenStudy (jdoe0001):

did you mean the "multiplicative inverse"?

OpenStudy (anonymous):

Yes

OpenStudy (jdoe0001):

well... can you make that mixed rational, to just an unmixed rational? :)

OpenStudy (jdoe0001):

what's -2 3/4 in just a fractional form?

OpenStudy (anonymous):

Can u just tell me the answer

OpenStudy (jdoe0001):

well, the kills the idea behind the exercise otherwise you'd have been handed out the answer sheets too, if all that's required is to transcribe them

OpenStudy (anonymous):

Can show me examples how to do it

OpenStudy (jdoe0001):

ok... lemme change the mixed fraction, to so-called improper fraction \(\bf -2\cfrac{3}{4}\implies \cfrac{-2\times 4+3}{4}\implies \cfrac{-5}{4}\\ \quad \\ \textit{multiplicative inverse of that is}\qquad \cfrac{4}{-5}\qquad why?\\ \quad \\ \cfrac{-5}{4}\times \cfrac{4}{-5}=1 \) is really just the same fraction upside-down

OpenStudy (jdoe0001):

hmmm.... hold one... one sec. . there... I have a mistake

OpenStudy (jdoe0001):

the negative is for the whole fraction... so.. one sec

OpenStudy (jdoe0001):

\(\bf -2\cfrac{3}{4}\implies -\cfrac{2\times 4+3}{4}\implies \cfrac{-11}{4}\\ \quad \\ \textit{multiplicative inverse of that is}\qquad \cfrac{4}{-11}\qquad why?\\ \quad \\ \cfrac{-11}{4}\times \cfrac{4}{-11}=1\)